# CLARITY: Contextual Linguistic Adaptation and Accent Retrieval for Dual-Bias Mitigation in Text-to-Speech Generation

### v1

@@ A.3.3 Bias Analysis
The P-value  [^12] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.binomtest.html of US and CA accents are shown in Table 6. The P-value indicates the result of the Binomial test, where a very small P-value (typically $p\leq 0.05$) means the observed proportion is significantly higher than the random chance, thus confirming the existence of a statistically significant bias. We can see that CosyVoice2 has a extreme and consistent bias towards US/CA accents across all tested non-US/CA accents. CLARITY_GPT demonstrates a much lower and more variable bias compared to CosyVoice2 which reduced overall bias. For British (GB), India (IN), Japan (JP), Malaysian (MY) and Singapore (SG) accents, there is no statistically significant evidence of bias toward US/CA accents. Although the magnitude of the bias is lower, CLARITY_GPT still exhibits a statistically significant bias for several accents, particularly Chinese (CN), Spanish (ES), Korean (KR) and Portuguese (PT). Bias towards US accent is more significant than CA accent.
Ablation Study: as shown in Figure 15, this ablation study details the impact of various model configurations on accent accuracy and systemic bias, demonstrating a phased improvement across experimental setups. The initial phase, which involved replacing the near-silent prompt audio with accent pool sourced prompt audio dynamically selected via instruction extracted speaker features, alone yielded a significant boost in both accent accuracy and fairness metrics. Building upon this foundation, performance saw further incremental gains through the subsequent integration of target text (standard) to prompt transcription similarity and the strategic use of prompt speech exhibiting a more pronounced accent (higher accent score). As mentioned in Section A.3.1, we also compared the results of standard and adapted text similarity. Ultimately, the most effective configurations (proposed) were achieved by leveraging Llama or GPT to perform adaptation on the target text, resulting in the highest overall accuracy and the lowest observed degree of bias among all tested methodologies.

### v2

@@ A-C3 Bias Analysis
The P-value  [^12] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.binomtest.html of US and CA accents are shown in Table VII. The P-value indicates the result of the Binomial test, where a very small P-value (typically $p\leq 0.05$) means the observed proportion is significantly higher than the random chance, thus confirming the existence of a statistically significant bias. We can see that CosyVoice2 has a extreme and consistent bias towards US/CA accents across all tested non-US/CA accents. CLARITY_GPT demonstrates a much lower and more variable bias compared to CosyVoice2 which reduced overall bias. For British (GB), India (IN), Japan (JP), Malaysian (MY) and Singapore (SG) accents, there is no statistically significant evidence of bias toward US/CA accents. Although the magnitude of the bias is lower, CLARITY_GPT still exhibits a statistically significant bias for several accents, particularly Chinese (CN), Spanish (ES), Korean (KR) and Portuguese (PT). Bias towards US accent is more significant than CA accent.
Ablation Study
As shown in Figure 15, this ablation study details the impact of various model configurations on accent accuracy and systemic bias, demonstrating a phased improvement across experimental setups. The initial phase, which involved replacing the near-silent prompt audio with accent pool sourced prompt audio dynamically selected via instruction extracted speaker features, alone yielded a significant boost in both accent accuracy and fairness metrics. Building upon this foundation, performance saw further incremental gains through the subsequent integration of target text (standard) to prompt transcription similarity and the strategic use of prompt speech exhibiting a more pronounced accent (higher accent score). As mentioned in Section A-C1, we also compared the results of standard and adapted text similarity. Ultimately, the most effective configurations (proposed) were achieved by leveraging Llama or GPT to perform adaptation on the target text, resulting in the highest overall accuracy and the lowest observed degree of bias among all tested methodologies.